1, A simple pendulum 0.4m long as a period of 1.3sec. what is the period of a similar pendulum 0.3m long in the same location.

1
Expert's answer
2022-01-10T09:06:45-0500

"T=2\pi\sqrt{\frac lg},"

"T\sim \sqrt l,"

"T_2=T_1\sqrt{\frac{l_2}{l_1}}=1.1~s."